OpenStudy (anonymous):

Jack's pet dog has been prescribed x mg of a medicine per day. Each pill of the medicine has 15 mg of the medicine. Jack has to give his dog a total of 25 pills. If Jack gave his dog medicine for 7 days, use the expression below to interpret the situation.

OpenStudy (anonymous):

the expression is 25 - 7x/15 Part A: What does x/15 represent? (3 points) Part B: What does the second term of the expression represent? (3 points) Part C: What does the entire expression represent

OpenStudy (triciaal):

1 pill = 15 mg ratio divide by 15 1/15 pill = 1mg x/15pill = x mg amount per day 7x/15 = amount for 7 days the total expression would represent the amount of pill over the time period